Congratulation! Probably you are right, but I could not get this is a real lock or something else.
1) How much was the fringe amplitude (DC) of the reflected beam? (Vref_max=XXX [V] and Vref_min=YYY [V])
Does this agree with the expectation?
2) Do you have the time series? (V_ref and V_error)
I guess I succeeded in locking of the cavity with the green beam
Strictly speaking, the laser frequency of the end NPRO is locked to the 40 meter arm cavity.
Pictures, some more quantitative numbers and some plots are going to be posted later.
After the alignment of the cavity I could see DC fringes in its reflection. Also I could see the cavity flashing on the monitor of ETMY_CCD.
I drove the pzt of the NPRO with f=200kHz, and then the spectrum analyzer showed 200kHz beat note in the reflection signal. This means it's ready to PDH technique.
And then I made a servo loop with two SR560s, one for a filter and the other for a sum amp.
After playing with the value of the gain and the sign of the feedback signal, the laser successfully got lock.
To make sure it is really locked, I measured the open loop transfer function of the PDH servo while it stayed locked. The result is shown in the attached figure.
The measured data almost agrees with the expected curve below 1kHz, so I conclude it is really locked.
However the plot looks very noisy because I could not inject a big excitation signal into the loop. If I put a big excitation, the servo was unlocked.
The current servo is obviously too naive and it only has f-1 shape, so the filter should be replaced by a dedicated PDH box as we planed.